OpenStudy (anonymous):

(sinx/1+cosx) + (sinx/1-cosx). i need help starting this to solve?):

jimthompson5910 (jim_thompson5910):

\[\Large \frac{\sin(x)}{1+\cos(x)} + \frac{\sin(x)}{1-\cos(x)}\] \[\Large \frac{\sin(x)(1-\cos(x))}{(1+\cos(x))(1-\cos(x))} + \frac{\sin(x)}{1-\cos(x)}\] \[\Large \frac{\sin(x)(1-\cos(x))}{(1+\cos(x))(1-\cos(x))} + \frac{\sin(x)(1+\cos(x))}{(1+\cos(x))(1-\cos(x))}\] does that help?

OpenStudy (anonymous):

yes:) then i got \[\frac{\sin^2x(1-cosx)(1+cosx) }{(1-cosx)(1+cosx) } \] would it just end up being \[\sin^2x\]

jimthompson5910 (jim_thompson5910):

close, but you would do this \[\Large \frac{\sin(x)}{1+\cos(x)} + \frac{\sin(x)}{1-\cos(x)}\] \[\Large \frac{\sin(x)(1-\cos(x))}{(1+\cos(x))(1-\cos(x))} + \frac{\sin(x)}{1-\cos(x)}\] \[\Large \frac{\sin(x)(1-\cos(x))}{(1+\cos(x))(1-\cos(x))} + \frac{\sin(x)(1+\cos(x))}{(1+\cos(x))(1-\cos(x))}\] \[\Large \frac{\sin(x)(1-\cos(x))+\sin(x)(1+\cos(x))}{(1+\cos(x))(1-\cos(x))}\] \[\Large \frac{\sin(x)-\sin(x)\cos(x)+\sin(x)+\sin(x)\cos(x)}{(1+\cos(x))(1-\cos(x))}\] \[\Large \frac{2\sin(x)}{(1+\cos(x))(1-\cos(x))}\] \[\Large \frac{2\sin(x)}{1-\cos^2(x)}\] \[\Large \frac{2\sin(x)}{\sin^2(x)}\] \[\Large \frac{2}{\sin(x)}\] \[\Large 2\csc(x)\]

OpenStudy (anonymous):

Oh! I had just cancelled the top and bottom two factors. Thank you so much! it really helped me!:)

jimthompson5910 (jim_thompson5910):

glad it did